Where does the -16 come from in the vertical motion equation?
Answer by Alan3354(69443)

Where does the -16 come from in the vertical motion equation?
----------
Earth's gravitational acceleration is apx 32.16 ft/sec/sec, 32 is often used.
16 is 1/2 of that, in the equation s = at^2/2 = 16t^2
It's negative because it's downward.
--> -16t^2